# Walking the Tightrope of LLMs for Software Development: A Practitioners’ Perspective

### v1

@@ Abstract
Background: Large Language Models emerged with the potential of provoking a revolution in software development (e.g., automating processes, workforce transformation). Although studies have started to investigate the perceived impact of LLMs for software development, there is a need for empirical studies to comprehend how to balance forward and backward effects of using LLMs. Objective: We investigated how LLMs impact software development and how to manage the impact from a software developer’s perspective. Method: We conducted 22 interviews with software practitioners across 3 rounds of data collection and analysis, between October (2024) and September (2025). We employed socio-technical grounded theory (STGT) for data analysis to rigorously analyse interview participants’ responses. Results: We identified the benefits (e.g., maintain software development flow, improve developers’ mental model, and foster entrepreneurship) and disadvantages (e.g., negative impact on developers’ personality and damage to developers’ reputation) of using LLMs at individual, team, organisation, and society levels; as well as best practices on how to adopt LLMs. Conclusion: Critically, we present the trade-offs that software practitioners, teams, and organisations face in working with LLMs. Our findings are particularly useful for software team leaders and IT managers to assess the viability of LLMs within their specific context.
@@ 1 Introduction
Many investigations (e.g., [14, 15, 16, 17]) highlight the potential benefits associated with software practitioners adopting LLMs. For instance, Cui et al. [15] found software developers using GitHub Copilot achieving an increase of 26.08% in the number of weekly completed tasks when conducting an experiment with almost five thousand software developers from companies including Microsoft and Accenture. On the other hand, many studies (e.g., [18, 19, 20, 21]) present the downsides of using LLMs. For instance, Lee et al. [20] observed from surveying 319 knowledge workers that LLMs can potentially affect software developers’ critical thinking skills.
@@
  • RQ1. How do LLMs take software developers forward? – Taking developers forward involves benefits gained from using LLMs.
  • RQ2. How do LLMs hold software developers back? – Holding developers back involves disadvantages arising from using LLMs.
  • RQ3. How do software developers achieve a balanced use of LLMs?
To answer these RQs, we conducted twenty-two semi-structured interviews with software practitioners across three rounds - between October (2024) and September (2025). Our study aims to comprehend the current industrial perspective on software developers adopting Large Language Model-based tools on SE-related activities, which involves exploring the benefits, challenges, limitations, and recommendations shared by software practitioners involved in SE-related activities. Our analysis reveals the following main benefits: (B1) reduced effort due to LLMs as a foundation to boost code development and perceived saving time; (B2) flow experience when LLMs mitigate interruptions and automate simple and tedious tasks; (B3) interaction with LLMs influencing developers’ personality and creating a safe space; and (B4) LLMs promoting LLM entrepreneurship as a consultant for every (not complex) question; and these disadvantages: (D1) disruption of the development flow due to an increase in developers’ effort influenced by LLMs; (D2) reduced effort due to LLMs negatively impacting developers’ personality (e.g., laziness) and hindering developers’ skills; (D3) interaction with LLMs reducing mentorship opportunities. The main contributions of this research include:
  • Identification and categorisation of benefits and disadvantages of adopting LLMs for software development tasks in terms of individual, team, organisation, and society level.
@@ 2 Methodology
Figure 1: Study methodology.
This study aims to comprehend the current industrial perspective on the impact of software practitioners adopting Large Language Model-based tools for SE-related activities. By employing the socio-technical grounded theory (STGT) for data analysis [25], we conducted three rounds of data collection and analysis involving interviews [^1] Approved by Monash Human Research Ethics Committee (No. 44875) with 22 software practitioners. Figure 1 provides an overview of our study methodology. In our first round, we conducted six interviews with novice software practitioners. In our second round, we conducted interviews with thirteen novice and experienced software practitioners, resulting in the emergence of concepts and categories. Finally, we interviewed three more software practitioners during the third round of data collection and analysis, with the aim of consolidating our findings. The supplementary material includes the pre-interview questionnaire, interview guide, STGT example, assumption list, and LLM capabilities, and is available online [26].
2.1 Why Socio-Technical Grounded Theory (STGT)?
In the face of many other qualitative research methods (e.g., thematic analysis, content analysis), we choose STGT, a variation of Grounded Theory (GT) [25, 27]. While STGT also includes human and social aspects similar to GT, it also acknowledges the importance of technical knowledge - a very relevant aspect of software engineering research - enabling more profound insights. In fact, STGT stands out as a research method suitable for topics where practice-based or industry perspectives are relevant.
To better explain the importance of using STGT in our context, we will explain our research topics using the four dimensions of its underlying socio-technical research framework:
  • Socio-technical Phenomenon: Our research investigates software practitioners’ perspectives on the role of LLM tools in software development. This phenomenon is characterised as socio-technical because the potential influence of individual aspects (e.g., motivations, emotions), team aspects (e.g., collaboration), and technology aspects (e.g., usability, features) can interplay on the experiences of software practitioners using LLM-powered tools.
  • Socio-technical Domain and Actors: The domain of LLMs for SE (LLM4SE), an intersection between Software Engineering and Artificial Intelligence, is a socio-technical domain since there is a “tight coupling between its social and technical aspects” [25, 27, 28]. The actors of this domain encompass software practitioners using LLM tools in software development, such as software developers, software engineers, data engineers, data scientists, and DevOps engineers.
  • Socio-technical researchers: This research was conducted by combining different experiences and skills among the research team. The interviews were conducted by an early-career researcher under the supervision of three experienced researchers.
  • Socio-technical data, tools, and techniques: During the study, the first author utilised Qualtrics survey to collect participant demographics, Zoom for recording and transcription, DeepL222www.deepl.com for translation of the interview transcriptions in Portuguese (P1, P4, P6) to English, and Nvivo and spreadsheets to support the analysis of the interview transcriptions.
Before starting the STGT study, it is fundamental to define how the STGT study will be conducted in terms of ontology, epistemology, and research paradigm. STGT offers flexibility in comparison with the traditional GT methods (i.e., Strauss-Corbinian GT, Glaserian GT, and Constructivist GT) with specific paradigms. Ontology is described as “What we believe exists or what we perceive as reality, in a research context” [27, Chapter 5]. Our study involves participants combining physical contexts and interactions (e.g., developers collaborating in a team physically) and virtual contexts and interactions (e.g., developers working remotely). Epistemology is described as “What can be treated as knowledge and how that knowledge is gained, in a research context” [27, Chapter 5]. Since we understand that the perceptions about benefits, challenges, and recommendations involving LLM adoption are subject to interpretation, we decided to adopt a subjective epistemology approach. Research paradigm is described as “Researcher worldview about what they believe is reality (ontology) and how knowledge about that reality can be gained (epistemology), in a research context” [27, Chapter 5]. Since we are adopting a subjective epistemology approach, believing in a socially constructed reality, we decided to follow a constructivist research paradigm.
@@ 2.2 Study Design and Piloting
Based on the findings (i.e., benefits, challenges, recommendations) from our systematic literature review [24], we developed a preliminary interview guide, which also included potential follow-up questions. This interview guide [26] was improved based on discussions with the PhD supervisors, as well as feedback from an industrial collaborator, who has experience managing a software team in an Australian company. The Attitude Towards Artificial Intelligence (ATAI) scale, developed by [29], was included to collect participants’ attitudes towards LLMs (AI). ATAI scale includes the following five nine-point scale questions, which we adapted to a five-point scale ranging from strongly disagree to strongly agree to facilitate the participants to answer them:
@@
Less than 2 years   7 Less than 1 year   8 3 - 5   5
Less than 1 year   5 Less than 2 years   5 11 - 20; 21 - 30   4
Less than 18 months   5 Less than 6 months   4 6 - 10   3
More than 2 years   5 Less than 18 months   3 Less than or equal to 2   2
@@
2.3 Sampling
STGT supports different sampling methods: purposive, random and convenience. Using our findings from [24] as a foundation, we began with purposive sampling in the first round, focusing on less experienced developers with less than five years of professional experience (See Fig. 1). In the following rounds, we moved to convenience sampling, interviewing software practitioners from all levels of experience. This would support further comparison between novice and experienced developers’ perceptions and experiences. We advertised this study on our professional social media, LinkedIn and X (formerly Twitter). We also recruited participants from our personal connections. Initially, we advertised our study, including a \$50 (AUD) voucher gift card. However, due to potential fake participants identified based on discrepancies between their IP addresses and the countries they submitted in the pre-interview questionnaire, we got approval from the Faculty Ethical Review Committee to omit the gift card in the advertisement, and only offered it to genuine participants at the end of the interview. We adopted a snowballing approach by encouraging the participants to share our study with others.
@@ 2.3 Sampling
Regarding their experience using LLM tools, which can influence the faced challenges and suggested solution strategies, participants reported using a variety of LLM tools. Not surprisingly, ChatGPT (n=18) was the most recurrent tool used by the interview participants. Figure II summarises participants’ demographics information related to the adoption of LLM tools. Most of the participants (n=14) have used LLMs in general, not specifically for SE tasks, for more than one year. According to their self-report amount of daily prompts, 59% (n=13) of participants engage daily with LLMs through at least 6 prompts. Their attitudes towards LLMs captured via Attitude Towards Artificial Intelligence (ATAI) scale highlight different inclinations towards AI technologies. While mostly (n=14) strongly disagreeing or somewhat disagreeing about fearing AI, they also mostly (n=14) disagree or somewhat disagree about trusting AI. Most (n=18) somewhat agree or strongly agree that AI will benefit humanity, and somewhat disagree or strongly disagree that AI will destroy humanity (n=14). However, most participants (n=13) somewhat agree or strongly agree that AI will affect the job market, leading to job losses. With respect to their experience using LLMs for SE tasks, mostly (n=16) reported having more than six months of experience, and they use LLMs mostly for code-related tasks (n=17), such as coding and debugging. Most of the study participants (n=14) reported that their companies allow them to employ LLM tools for job responsibilities.
The fast-paced evolution of LLM tools compels us to examine the features available in the tools mentioned by interview participants. During our three rounds of data collection and analysis, we observed an emergence and evolution of different LLM tools (e.g., ChatGPT, Cursor) and integration with traditional tools (e.g., ChatGPT for Databricks [30], Copilot for Power BI [31]). This was also highlighted by the participants, e.g.: 9 “I’ve been able to see a bit of how those tools evolve during this time. [For example, GitHub] Copilot evolved from being a better autocomplete to having more tools [and features].” – P10 [Software Engineer]. Ferino et al. [24] suggest researchers investigating LLMs for SE to provide a snapshot of the LLM features with the intent of improving clarity of the context related to the data collection. We summarise the main features in the supplementary online package [26].
Figure 2: Emergence of the category “Impact on using LLMs” from raw datacodesconceptssubcategoriescategory through constant comparison.
@@ 2.4 Data Analysis
One of the ways in which STGT distinguishes itself from traditional grounded theory methods is by being organised into two stages: basic stage of data collection and analysis, and advanced stage for theory development. The basic stage of data collection analysis involves conducting a lean literature review to identify research gaps, followed by data collection (e.g., survey), and open coding, constant comparison, and basic memoing for data analysis. When managed carefully, systematic literature reviews (SLRs) [24] can also be conducted. The advanced stage of theory development involves conducting targeted data collection and analysis, aiming to reach theoretical saturation, theory structuring, and a targeted literature review. Through this clear separation, STGT gives flexibility to the researcher to decide whether to proceed to the advanced stage or share and communicate the findings from the basic stage [28]. Driven by the rich and fascinating results from the basic stage, we decided to communicate the findings.
@@
Open Coding. The process of open coding was iteratively improved by the first researcher following the guidance of an STGT expert, the second author. Initially, the first author printed out one of the interviews and coded it, followed by a review and discussions with an experienced researcher. After this, the researcher coded two more interviews using Google Docs. The researcher decided to move to NVivo, instigated by its features, and spreadsheets. Before doing the open coding, and between rounds of data collection and analysis, the first author was advised to create and update an assumption list [26]. This helped to make transparent the researcher’s bias.
@@
Constant Comparison. By doing this, similar codes were grouped into concepts, and similar concepts under sub-categories, and similar sub-categories under categories. During this process, we drew diagrams to facilitate visualisation and insights regarding the relationship between concepts and sub-categories, and also conducted discussions between the authors. Figure 2 illustrates the emergence of the category Impact of using LLMs via constant comparison.
@@
Memoing. While performing constant comparison, we also wrote down memos reflecting our ideas, thoughts, and key elements emerging throughout the analysis. Those memos, which encompass both interview and inter-interview levels, supported the update of the interview guide between rounds. At the interview level, we compared the participants’ pre-interview questionnaire responses with their interview responses, especially concerning the ATAI scale. That information provided more context for our analysis. At the inter-interview level, we compared codes from different interview participants, which also helped to refine our concepts and sub-categories. An example of a memo “How LLMs affect developers’ intuition?” is shown below. The discussion on the main insights from memoing is presented in Section 4.2.
@@
Memo on “How LLMs affect developers’ intuition?” LLMs seem to emerge as a learning tool for software developers based on P12, P14, and P15, e.g., relying on LLMs to cover gaps in knowledge. At the same time, P14 argues that while using LLMs for learning, novice developers should do the majority of the work. This practice seems to be necessary to improve novice developers’ skills, but also to cultivate the intuition inherent in experienced developers. Intuition is something that comes naturally after fully understanding something. That intuition (internalised knowledge) helps to avoid scenarios where LLMs are misleading, which P11 described as ChatGPT being a very good liar.
@@ 3 Findings
Tables 3.1.1 and 3.2.1 illustrate the benefits and disadvantages at the individual level of using LLM tools to support software development tasks, along with the reason influencing these impacts. For instance, when using LLMs for brainstorming, it has a negative impact on developers’ personalities because of overreliance on LLMs, reduction of effort, mitigation of interruptions in software development and also because of the loss of control over LLMs. Additionally, we have summarised the main benefits and disadvantages in Figure 3 and 4, respectively. For instance, using LLMs for debugging can boost code development, saving development time, reduce software developers’ effort, and provide learning opportunities to software developers. We organised the remainder of this section by describing benefits in using LLMs (section 3.1), disadvantages in using LLMs (section 3.2), and recommendations on how to use LLMs (section 3.3). Throughout this section, we provide representative quotes to illustrate participants’ experiences and viewpoints. In order to protect the privacy of our participants, we refer to the interview participants as P1-P22. Similar to Masood et al.[32], we will use some expressions to indicate the extent of the predominance. In this sense, ‘few’ refers to less than 25%, ‘many or majority’ to over 50%. Although this is not a quantitative study, it indicates how much evidence supports each theme.
3.1 RQ1. How do LLMs take software developers forward?

### v2

@@ Abstract
Background: Large Language Models emerged with the potential of provoking a revolution in software development (e.g., automating processes, workforce transformation). Although studies have started to investigate the perceived impact of LLMs for software development, there is a need for empirical studies to comprehend how to balance forward and backward effects of using LLMs. Objective: We investigated how LLMs impact software development and how to manage the impact from a software developer’s perspective. Method: We conducted 22 interviews with software practitioners across 3 rounds of data collection and analysis, between October (2024) and September (2025). We employed Socio-Technical Grounded Theory for Data Analysis (STGT4DA) to rigorously analyse interview participants’ responses. Results: We identified the benefits (e.g., maintain developer flow, improve developer mental models, and foster entrepreneurship) and challenges (e.g., damage to developers’ reputation) of using LLMs at individual, team, organisation, and society levels; as well as actionable guidances into how mitigate these challenges. Conclusion: Critically, we present the trade-offs that software practitioners, teams, and organisations face in working with LLMs. Our findings are particularly useful for software team leaders and IT managers to assess the viability of LLMs within their specific context.
@@ 1 Introduction
Many investigations (e.g., [59, 9, 16, 43]) highlight the potential benefits associated with software practitioners adopting LLMs. For instance, Cui et al. [9] found that software developers using GitHub Copilot achieved an increase of 26.08% in the number of weekly completed tasks when conducting an experiment with almost five thousand software developers from companies including Microsoft and Accenture. On the other hand, many studies (e.g., [31, 30, 32, 8]) present the downsides of using LLMs. For instance, Lee et al. [32] observed from surveying 319 knowledge workers that LLMs can potentially affect software developers’ critical thinking skills.
@@
  • RQ1. How do LLMs help software developers move forward? – Moving developers forward involves the experienced benefits of using LLMs.
  • RQ2. How do LLMs create an imbalance in software development? – Creating an imbalance involves experienced challenges when using LLMs. We draw special attention to this RQ in this report, as its findings may guide future studies.
  • RQ3. How can software developers achieve a balanced use of LLMs? – It encompasses actionable recommendations to use LLMs.
To answer these RQs, especially RQ2, we conducted twenty-two semi-structured interviews with software practitioners across three rounds - between October (2024) and September (2025). Our study aims to comprehend the current industrial perspective on software developers adopting Large Language Model-based tools on SE-related activities, which involves exploring the benefits, challenges, limitations, and recommendations shared by software practitioners involved in SE-related activities. Our analysis reveals the following main benefits: (B1) reduced effort due to LLMs as a foundation to boost code development and perceived saving time; (B2) flow experience when LLMs mitigate interruptions and automate simple and tedious tasks; (B3) interaction with LLMs influencing developers’ personality and creating a safe space; and (B4) LLMs promoting LLM entrepreneurship as a consultant for every (not complex) question; and these challenges: (D1) developer flow interruptions due to frictions in human-LLM interaction; (D2) reduced effort due to LLMs eroding developer skills and agency; (D3) interaction with LLMs reducing mentorship opportunities. The main contributions of this research include:
  • Identification and categorisation of benefits and challenges of adopting LLMs for software development tasks in terms of individual, team, organisation, and society level.
@@ 2 Methodology
Figure 1: Iterative steps of data collection and analysis.
This study aims to comprehend the current industrial perspective on the impact of software practitioners adopting LLM-based tools for SE-related activities. We conducted an interview-based study with 22 software practitioners across three rounds of data collection [^1] Approved by Monash Human Research Ethics Committee (No. 44875). We applied socio-technical grounded theory for data analysis (STGT4DA) [22] to analyse the interview data iteratively. In our first round, we conducted six interviews with novice software practitioners. In our second round, we conducted interviews with thirteen novice (n=10) and experienced (n=3) software practitioners, resulting in the emergence of concepts and categories. Finally, we interviewed three more software practitioners during the third round of data collection and analysis, with the aim of consolidating our findings. Figure 1 provides an overview of our study methodology. The supplementary material, which includes the pre-interview questionnaire, interview guide, STGT4DA coding example, assumption list, and LLM capabilities, is available online [18].
@@ 2.1 Study Design and Piloting
Based on the findings (i.e., benefits, challenges, recommendations) from our systematic literature review [19], we developed a preliminary interview guide, which also included potential follow-up questions. This interview guide [18] was improved based on discussions with the PhD supervisors, as well as feedback from an industrial collaborator, who has experience managing a software team in an Australian company. The Attitude Towards Artificial Intelligence (ATAI) scale, developed by Sindermann et al. [51], was included in the pre-interview questionnaire to provide an initial understanding of participants’ attitude towards AI, allowing the interviewer (first author) to adjust interview questions based on each participant’s strongest responses across five dimensions — perceived benefits, trust, fear, job displacement, and existential concerns — and also afforded additional contextual information to support the subsequent data analysis (i.e., open coding and memoing). ATAI scale includes the following five nine-point scale questions, which we adapted to a five-point scale ranging from strongly disagree to strongly agree to facilitate the participants to answer them:
@@
2.2 Sampling
Leveraging the findings from our SLR [19] on novice developers’ adoption of LLMs for SE tasks as a foundation, we began with purposive sampling in the first round, focusing on less experienced developers with less than five years of professional experience (See Fig. 1). In the following rounds, we moved to convenience sampling, interviewing software practitioners from all levels of experience since our first round of data collection and analysis did not result in enough codes, concepts, or memos to justify continuing only with novice participants. This would also support further comparison between novice and experienced developers’ perceptions and experiences. We advertised this study on our professional social media, LinkedIn and X (formerly Twitter). We also recruited participants from our personal connections. Initially, we advertised our study, including a \$50 (AUD) voucher gift card. However, due to potential fake participants identified based on discrepancies between their IP addresses and the countries they submitted in the pre-interview questionnaire, we got approval from our Faculty Ethical Review Committee to omit the gift card in the advertisement, and only offered it to genuine participants at the end of the interview. We also adopted a snowballing approach by encouraging the participants to share our study with others.
@@ 2.2 Sampling
Less than 1 year   5 Less than 6 months   4 Less than or equal to 2   2
Less than 1.5 year   5 Less than 1 year   8 3 - 5   5
Less than 2 years   7 Less than 1.5 year   3 6 - 10   3
More than 2 years   5 Less than 2 years   5 11 - 20; 21 - 30   4
@@
Regarding their experience using LLM tools, which can influence the faced challenges and suggested solution strategies, participants reported using a variety of LLM tools. Not surprisingly, ChatGPT (n=18) was the most recurrent tool used by the interview participants. Table II summarises participants’ demographics information related to the adoption of LLM tools. Most of the participants (n=14) have used LLMs in general, not specifically for SE tasks, for more than one year. According to their self-report amount of daily prompts, 59% (n=13) of participants engage daily with LLMs through at least 6 prompts. Their attitudes towards LLMs captured via the Attitude Towards Artificial Intelligence (ATAI) scale highlight different inclinations towards AI technologies. While mostly (n=14) strongly disagreeing or somewhat disagreeing about fearing AI, they also mostly (n=14) disagree or somewhat disagree about trusting AI. Most (n=18) somewhat agree or strongly agree that AI will benefit humanity, and somewhat disagree or strongly disagree that AI will destroy humanity (n=14). However, most participants (n=13) somewhat agree or strongly agree that AI will affect the job market, leading to job losses. With respect to their experience using LLMs for SE tasks, mostly (n=16) reported having more than six months of experience, and they use LLMs mostly for code-related tasks (n=17), such as coding and debugging. Most of the study participants (n=14) reported that their companies allow them to employ LLM tools for job responsibilities.
The fast-paced evolution of LLM tools compels us to examine the features available in the tools mentioned by interview participants. During our three rounds of data collection and analysis, we observed an emergence and evolution of different LLM tools (e.g., ChatGPT, Cursor) and integration with traditional tools (e.g., ChatGPT for Databricks [14], Copilot for Power BI [54]). This was also highlighted by the participants, e.g.: “I’ve been able to see a bit of how those tools evolve during this time. [For example, GitHub] Copilot evolved from being a better autocomplete to having more tools [and features].” – P10 [Software Engineer]. Ferino et al. [19] suggest that researchers investigating LLMs for SE should provide a snapshot of the LLM features with the intent of improving clarity of the context related to the data collection. We summarise the main features in the supplementary online package [18].
Figure 2: Emergence of the category “Impact on using LLMs” from raw datacodesconceptssubcategoriescategory through constant comparison [23].
@@ 2.3 Data Analysis
Applying STGT’s basic stage of data collection and analysis, we conducted a lean literature review to identify research gaps, followed by data collection (e.g., survey), and hashtag open coding, constant comparison, and basic memoing for data analysis [22, 23]. STGT4DA has been applied in various SE studies to identify “important patterns in the qualitative data and present them as layered and/or multi-dimensional findings, along with insights and reflections” [46, 36, 37, 21]. We leveraged its foundational socio-technical research framework to investigate the socio-technical phenomenon of software practitioners’ perspectives on the role of LLM tools in software development, in the socio-technical domain of LLMs for SE (LLM4SE), with socio-technical actors such as software developers, software engineers, data engineers, data scientists, and DevOps engineers. As socio-technical researchers, we combined our skills and experiences. The interviews were conducted by an early-career researcher (first author) under the guidance of three experienced researchers, and the data were analysed by the first author with detailed feedback from the second author. We also applied socio-technical data, tools, and techniques such as a Qualtrics survey to collect participant demographics, Zoom for recording and transcription, DeepL222www.deepl.com for translation of the interview transcriptions in Portuguese (P1, P4, P6) to English, and Nvivo and spreadsheets to support the analysis of the interview transcriptions.
@@
Hashtag Open Coding. The process of hashtag open coding [23, Chapter 10] was iteratively improved by the first researcher following the guidance of the second author. Initially, the first author printed out one of the interviews and coded it, followed by a review and discussions with an experienced researcher. After this, the researcher coded two more interviews using Google Docs. The researcher decided to move to NVivo, instigated by its features, and spreadsheets. Before performing the hashtag open coding, and between rounds of data collection and analysis, the first author was advised to create and update an assumption list [18]. This helped to make the researcher’s assumptions and biases transparent so they could be managed.
@@
Constant Comparison. Similar codes were grouped into concepts, and similar concepts under sub-categories, and similar sub-categories under categories. During this process, we drew diagrams to facilitate visualisation and insights regarding the relationship between concepts and sub-categories, and also conducted discussions between the authors. Figure 2 illustrates the emergence of the category Impact of using LLMs via coding and constant comparison.
@@
Memoing. While performing constant comparison, we also wrote down memos reflecting our ideas, thoughts, and key elements emerging throughout the analysis. Those memos, which encompass both interview and inter-interview levels, supported the update of the interview guide between rounds. At the interview level, we compared the participants’ pre-interview questionnaire responses with their interview responses, especially concerning the ATAI scale. That information provided more context for our analysis. At the inter-interview level, we compared codes from different interview participants, which also helped to refine our concepts and sub-categories. An example of a memo “How do LLMs affect developer intuition?”, is shown below. The discussion on the main insights from memoing is presented in Section 4.2.
@@
Memo on “How do LLMs affect developer intuition?” LLMs seem to emerge as a learning tool for software developers based on participants P12, P14, and P15, e.g., relying on LLMs to cover gaps in knowledge. At the same time, P14 argues that while using LLMs for learning, novice developers should do the majority of the work. This practice seems to be necessary to improve novice developers’ skills, but also to cultivate the intuition inherent in experienced developers. Intuition is something that comes naturally after fully understanding something. That intuition (internalised knowledge) helps to avoid scenarios where LLMs are misleading, which P11 described as ChatGPT being “a very good liar”.
@@ 3 Findings
Tables 3.1.1 and LABEL:tab:disadvantageIndividualLevel illustrate the benefits and challenges at the individual level of using LLM tools to support software development tasks, along with the reasons influencing these impacts. For instance, Table LABEL:tab:disadvantageIndividualLevel shows that using LLMs for brainstorming may hinder developers’ skills due to a lack of practice in software development skills and overreliance on LLMs. Additionally, we have summarised the main benefits and challenges in Figures 3 and LABEL:fig:MainDisadvantagesIndividualLevel, respectively. For instance, using LLMs for debugging can boost code development, saving development time, reducing software developers’ effort, and providing learning opportunities to software developers. We organised the remainder of this section by describing thebenefits of using LLMs (section 3.1), challenges when using LLMs (section 3.2), and recommendations on how to use LLMs (section LABEL:subsec:recommendations). Throughout this section, we provide representative quotes to illustrate participants’ experiences and viewpoints. In order to protect the privacy of our participants, we refer to the interview participants as P1-P22. Similar to Masood et al.[38], we will use some expressions to indicate the extent of the predominance. In this sense, ‘few’ refers to less than 25%, ‘many or majority’ to over 50%. Although this is not a quantitative study, it indicates how much evidence supports each theme.
3.1 RQ1. How do LLMs help software developers move forward?